OpenStudy (anonymous):

I need help (b) Write an inequality to show how many snacks of one type would need to be sold to earn a profit of at least $500. Support your inequality with an explanation. Info: The club sold the snacks for $3 each. The club gets half of this amount as profit. The club also gets a bonus of $100 for any snack when it sells at least 200 of that snack. Last year’s data on sales are shown in the table. Type of Snack Number sold Energy Bars: 194 Baked chips: 200 Organic treats: 235 Whole grain crackers: 78

OpenStudy (anonymous):

I'd really appreciate if anyone can take the time to read it, I just have to make an inequality expression and solve it

OpenStudy (anonymous):

Take all the time you need, not too long though ^_^

OpenStudy (anonymous):

I kind of got half way through: (more than and equal sign) 500 I don't get the first part though

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

Was there a part (a) and anything before this?

OpenStudy (anonymous):

Yes, (a) Which snack(s) earned a profit of at least $350? Show your work.

jimthompson5910 (jim_thompson5910):

ok so for part (b), if we made x the number of snacks sold, then how much money do we make if each snack is sold for $3 ?

OpenStudy (anonymous):

wait so would the expression be 3x (more than or equal sign) 500?

jimthompson5910 (jim_thompson5910):

close, there are a few missing pieces though

jimthompson5910 (jim_thompson5910):

`The club gets half of this amount as profit. ` so we take the `3x` and cut that in half to get `3x/2`

jimthompson5910 (jim_thompson5910):

now what is the value of `3x/2` when x = 200 ?

OpenStudy (anonymous):

300?

jimthompson5910 (jim_thompson5910):

yes, since that isn't 500 or more, we haven't reached the target but there's a $100 bonus when 200 snacks are sold but 300+100 = 400 is still not enough, which means you need to sell more snacks

jimthompson5910 (jim_thompson5910):

in the end, the inequality set up will be \[\Large \frac{3x}{2} + 100 \ge 500\]

OpenStudy (anonymous):

ah ok

OpenStudy (anonymous):

So that's all? If so, that was fast! Thanks

jimthompson5910 (jim_thompson5910):

yeah that's all there is to it 3x/2 is the profit made, then you add on the $100 bonus (it's added on because 200 snacks aren't enough to reach $500 profit, so more need to be sold)

OpenStudy (anonymous):

Ok, thanks for the explanation! Helped a lot!!

jimthompson5910 (jim_thompson5910):

no problem